OpenStudy (anonymous):

Write in exponential form logbase4(x-4)=2

OpenStudy (ash2326):

@Courtney.soden @myininaya will help you with this

myininaya (myininaya):

\[\log_{b}(x)=y \text{ is log form }\] \[b^y=x \text{ is exponential form }\]

myininaya (myininaya):

So knowing this what is b?

OpenStudy (anonymous):

b=4

myininaya (myininaya):

ok and what is y?

OpenStudy (anonymous):

y=2

myininaya (myininaya):

\[\log_b(a)=y => b^y=a \] (I'm changing that x to a so it looks less confusing ) So a is x-4 Just plug into the exponential form and boom you got it! :)

myininaya (myininaya):

So replacing b with 4 we have \[4^y=a\] Replacing y with 2 we have \[4^2=a\] And replace a with x-4 what do we have?

OpenStudy (anonymous):

4^2=x-4

myininaya (myininaya):

yep :)

OpenStudy (anonymous):

thank you